# Area Law for the entanglement entropy of free fermions in nonrandom ergodic field

### v2

@@ Abstract.
The paper deals with the asymptotic behavior of one of the widely used characteristics of correlations in large quantum systems. The correlations are known as quantum entanglement, the characteristic is called the entanglement entropy, and as large systems we consider an ideal gas of spinless lattice fermions. The system is determined by its one-body Hamiltonian. As shown in [18], if the Hamiltonian is an ergodic finite difference operator with exponentially decaying spectral projection, then the asymptotic form of the entanglement entropy is the so-called Area Law. However, the only one-body Hamiltonian for which this spectral condition is verified is the $d$-dimensional discrete Schrödinger operators with random potential. In the present paper, we prove that the same asymptotic form of the entanglement entropy holds for a wide class of Schrödinger operators whose potentials are ergodic but nonrandom. We start with the quasiperiodic and limit-periodic operators, and then pass to the interesting and highly non-trivial case of the potentials generated by subshifts of finite type. They arose in the theory of dynamical systems in the study of non-random chaotic phenomena. As it turns out, obtaining the asymptotics of the entanglement entropy of free fermions requires a quite involved spectral analysis of the corresponding Schrödinger operator. Specifically, we prove for this class two important and interesting in itself spectral properties, known as exponential dynamical localisation in expectation and the exponential decay of the eigenfunction correlator, implying the Area Law for the entanglement entropy.
@@ 1. Introduction
Nevertheless, the observation allows us to expect that the Area Law holds for the classes of ergodic operators other than random ones, provided that they also exhibit “strong” Anderson localisation. In particular, these are finite difference operators with quasi-periodic potentials, defined on an orbit of an irrational winding on the torus, or with potentials that are functions defined on an orbit of a more complex finite-dimensional dynamical systems, including the doubling map and the famous Arnold’s cat map [9, 1].
The goal of this paper is to prove the Area Law for these classes of ergodic operators. The paper is organised as follows. In Section 2 we present our results and comments on them. The section contains four theorems and certain remarks establishing the validity of the Area Law for free lattice fermions whose one-body operator is a finite-difference operator in $\ell^{2}(\mathbb{Z}^{d})$ with dynamically generated coefficients. Specifically, these are self-adjoint ergodic operators that are the sums of a convolution operator (1.13) (the discrete Laplacian in the case of Schrödinger operator (1.4)) and a dynamically generated potential, a quasi-periodic function in particular, see (2.1) – (2.2) and (2.3) below. Theorems 1 and 2 treat multidimensional and one-dimensional quasiperiodic and limit-periodic cases respectively. For all these cases the spectral theory is well enough developed. Thus, the proofs of the theorems are quite straightforward and based mostly on these developments completed by certain arguments of spectral theory and corresponding references.
Theorem 3 treats the case of the one-dimensional Schrödinger operators with potentials generated by a subshift of finite type, an interesting dynamical system that is more complicated than the irrational shift on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, generating quasi-periodic potentials. Here, the already existing nontrivial spectral theory must nevertheless be substantially extended. We develop a necessary extension and use it to prove the validity of the Area Law for the corresponding operators, as well as their certain spectral properties, including that known as the exponential dynamical localisation in expectation.
In addition, we obtain a useful criterion for the exponential decay of the eigenfunction correlator for a general bounded one-dimensional discrete Schrödinger operator in terms of control over the “bad” spectral parameters on large but distant boxes (see Lemma 4.4). We believe that this criterion is of independent interest. We use this criterion in the proof of Theorem 4, establishing the exponential decay of the eigenfunction correlator for the one-dimensional Schrödinger operators with potentials generated by a subshift of finite type.
@@ 2. Results and Comments
and its particular case – Schrödinger operator (1.4), where
$W(\mathbf{n})=\delta_{\mathbf{n}+\mathbf{e}_{j}}+\delta_{\mathbf{n}-\mathbf{e}_{j}},\;\mathbf{n}\in\mathbb{Z}^{d},\;j=1,\dots,d,$
@@ 2.1. Results
    with $\|\langle\mathbf{n},\mathbf{\alpha}\rangle\|_{d}=\mathrm{dist}(\langle\mathbf{n},\mathbf{\alpha}\rangle,\mathbb{Z})$ and the Fermi energy
    $\varepsilon_{F}\in\sigma_{pp}(H)=\sigma(H)=\mathbb{R}$
    ((cf. Theorem 2 ((i)) about the 1d case below)).
@@
$(i)$ There is a one-dimensional version of Theorem 1 $(iv)$. It was considered by Damanik and Gan in [13] where an explicit condition on the Cantor group $\Omega$ is formulated.
@@
    The one-dimensional Maryland model: the one-dimensional discrete Schrödinger operator (2.12) with the potential ((cf. (2.4)))
@@
    and with the Fermi energy (1.6) $\varepsilon_{F}\in\sigma_{-}(H_{\omega})$, where $\sigma_{-}(H_{\omega})$ is the lower part of the dense pure point spectrum of the model
    $\sigma_{pp}(H_{\omega})\supseteq\big\{\lambda\in\mathbb{R}:\gamma(\lambda,g)>\beta(\alpha)\big\}=\sigma_{-}(H_{\omega})\cup\sigma_{+}(H_{\omega}),$
    (2.15) $\sigma_{-}(H_{\omega})=(-\infty,-\varepsilon_{0}],\;\;\sigma_{+}(H_{\omega})=[\varepsilon_{0},\infty),\;\varepsilon_{0}\geq 0,$
    where $\beta(\alpha)$ is given in (2.13) and $\gamma(\lambda,g)\geq\gamma(0,g)>0$ is the Lyapunov exponent of the corresponding finite-difference equation of the second order (B.2) ((see (4.5) for the general definition of the Lyapunov exponent, and Appendix B for the form of $\gamma(\lambda,g)$ and $\varepsilon_{0}$ in the one-dimensional Maryland model)), and $\varepsilon_{0}=0$ if $\gamma(0,g)>\beta(\alpha)$.
@@
The proof of the theorem is given in the next section.
@@
Using recent results of Jitomirskaya and Kachkovskiy [27, Corollary 5.14], it is possible to obtain a stronger version of Theorem 2 $(ii)$ in the following more general setting. Assume that $v:\mathbb{R}\rightarrow[-\infty,+\infty)$ is a $1$-periodic function, satisfying ((cf. (2.7)))
$v(y)-v(x)\geq\gamma(y-x),\quad\text{for}\quad\gamma>0,\quad 0\leq x<y<1.$
In particular, $v$ is bounded on $(0,1)$, but not necessarily finite at x=0,1(x=0,1(see [27] for the analysis of possible discontinuities of $v$ at $0,1)$ . Assume that
$\int_{0}^{1}\log(1+|v(x)|)dx<+\infty.$
Consider the one-dimensional Schrödinger operator (2.12) with potential
$V(\omega,n)=v(\omega+n\alpha),\;\omega\in\mathbb{T},\,\alpha\in\mathbb{T},$
and $\varepsilon_{F}\in\sigma_{-}(H_{\omega})$, where $\sigma_{-}(H_{\omega})$ is the lower part of the dense pure point spectrum of the model ((cf. (2.15)))
$\sigma_{pp}(H_{\omega})\supseteq\big\{\lambda\in\mathbb{R}:\gamma(\lambda)>\beta(\alpha)+\delta\big\}=\sigma_{-}(H_{\omega})\cup\sigma_{+}(H_{\omega}),\delta>0,$
$\sigma_{-}(H_{\omega})=(-\infty,-\varepsilon_{0}],\;\;\sigma_{+}(H_{\omega})=[\varepsilon_{0},+\infty),\;\varepsilon_{0}\geq 0,$
where $\beta(\alpha)$ is given in (2.13) and $\gamma(\lambda)\geq\gamma(0)>0$ is the corresponding Lyapunov exponent $(\varepsilon_{0}=0$ if $\gamma(0)>\beta(\alpha))$.
The proof is the same as for Theorem 2 $(ii)$ using the ULE property (3.12) established in [27, Corollary 5.14].
In our next result, we consider one-dimensional discrete Schrödinger operators (2.12) with potentials generated by a subshift of finite type (see Section 4), a very interesting class of hyperbolic dynamical systems exhibiting the so-called classical dynamical chaos, see e.g. [9]. For these ergodic operators, the spectral localisation was recently proved in [6, 7].
@@
In the proof of Theorems 1 and 2 in Section 3 and Appendices AC we essentially use the existing results of spectral analysis of the operators in question. To prove Theorem 3 in Section 4 we have to substantially extend the existing spectral theory. We believe that this extension (see Theorem 4 below) is of independent interest for the spectral theory of this class of operators and completes the results of important recent works [7, 6].
@@
In these cases we define the potential by $V_{\omega}(n)=gv(T^{n}\omega)$ with $v:\Omega_{A}\rightarrow\mathbb{R}$ satisfying the assumptions of Theorem 3. Then, there exists $g_{0}>0$ such that for all $0<g\leq g_{0}$ Theorems 3 and 4 hold true. Moreover, as pointed out in [7, Remark 2.18], $g_{0}$ is not too small if $\alpha$ is not too small.
@@ 2.2. Comments and Related Results
The proof of Theorems 3 and 4 for the Schrödinger operators whose potential is generated by a subshift of finite type is using as an input the uniform positivity of the Lyapunov exponent and the uniform large deviation-type estimate from [6, 7]. In addition, we use special techniques of the theory of dynamical systems involving the Markov partitions and related Markov chains [1, 9].
@@ 3. Finite-difference operators with quasi-periodic potential
In this section, we present the proofs of Theorem 1 and Theorem 2 on the Area Law for several classes of non-random ergodic operators. To the best of our knowledge, the only class of these operators for which the Area Law is rigorously established (using Criterion 1) is that consisting of the Schrödinger operators with independent identically distributed random potentials (aka Anderson model) whose probability law possesses a certain amount of smoothness [18].
@@
that is, $\Omega=\mathbb{T}$ equipped with the Borel $\sigma$-algebra and $\mathbf{P}$ being the normalized Lebesgue measure.
@@
It is shown in Appendix A that
@@
where $C<\infty$ and $c>0$ are independent of $\lambda\in\mathbb{R}$ and $\mathbf{n}\in\mathbb{Z}^{d}$. Combining (3.3) and (3.4), we obtain
@@
where $C<\infty$ and $c>0$ are independent of $\omega$. This and spectral theorem for the Fermi projection

### v3

@@ Abstract.
This paper deals with the asymptotic behaviour of a widely used correlation characteristic in large quantum systems. The correlations are known as quantum entanglement, the characteristic is called entanglement entropy, and the system is an ideal gas of spinless lattice fermions. The system is determined by its one-body Hamiltonian. It is shown in EPS [18] that if the Hamiltonian is an ergodic finite difference operator with an exponentially decaying spectral projection, then the asymptotic form of the entanglement entropy is the so-called area law. However, the only class of one-body Hamiltonians for which this spectral condition was verified is that consisting of discrete Schrödinger operators with random potential. In this paper, we prove the validity of the area law for several classes of Schrödinger operators whose potentials are ergodic but not random. We begin with quasiperiodic and limit-periodic operators and then move on to the interesting and highly non-trivial case of potentials generated by subshifts of finite type. These arose in the theory of dynamical systems when studying non-random chaotic phenomena. The corresponding asymptotic study requires quite an involved spectral analysis. Consequently, the majority of the paper is devoted to the proof and application of a variety of spectral properties of the operators in question, in particular we prove uniform localisation of the ejgenfunctions for the Maryland model and the exponential decay of the eihgenfunction correlator for a variety of models . We believe that these properties are of considerable independent interest.
@@ 1. Introduction
Nevertheless, the observation allows us to expect that the Area Law holds for the classes of ergodic operators other than random ones, provided that they also exhibit sufficiently “strong” Anderson localisation. In particular, these are finite difference operators with quasi-periodic potentials, defined on an orbit of an irrational winding on the torus, or with potentials that are functions defined on an orbit of a more complex finite-dimensional dynamical systems, including the doubling map and the famous Arnold’s cat map [9, 1].
The goal of this paper is to prove the Area Law for these classes of ergodic operators. The paper is organised as follows. In Section 2 we present our results on the validity of the Area Law for free lattice fermions whose one-body Hamiltonian is a finite-difference operator in $\ell^{2}(\mathbb{Z}^{d})$ with dynamically generated coefficients. Specifically, these are self-adjoint ergodic operators that are the sums of a convolution operator (1.13) (the discrete Laplacian in the case of Schrödinger operator (1.4)) and a dynamically generated potential, see (2.1) – (2.2) and (2.4) below. Theorems 1 and 2 treat multidimensional and one-dimensional quasiperiodic and limit-periodic cases. Theorem 4 treats the case, where the one-body Hamiltonian is the one-dimensional Schrödinger operator with potentials generated by a subshift of finite type, a complex dynamical system that is much more complicated than the irrational shift on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, which generates quasi-periodic potentials.
All of these operators have a purely point component of spectrum. However, our proof, based on the results of [18], makes significant use of some form of exponential bounds on the spectral (Fermi) projection of the operator in question, i.e., a considerably stronger property. In certain cases this property is partially established. There are, however, several important cases where the existing spectral theory has to be considerably developed. The first is the case of the archetype Maryland model, both multi-dimensional and one-dimensional, where we have to prove the property of uniform localisation of eigenfunctions, presented in Theorem 3. The second important case is the Schrödinger operator whose potentials generated by the subshift of finite type. Here we have to prove the exponential decay of the eigenfunction correlator, presented in Theorem 5. This includes a useful criterion for the exponential decay of the eigenfunction correlator for ergodic bounded one-dimensional discrete Schrödinger operators in terms of control over the “bad” sets of spectral parameters of the operator restriction to large but distant boxes, see Lemma 2.5. We believe that these spectral results are of independent interest.
@@ 2. Results and Comments
and its particular case – Schrödinger operator (1.4), where $W$ is the $d$-dimensional discrete Laplacian
(2.3) $W(\mathbf{n})=\delta_{\mathbf{n}+\mathbf{e}_{j}}+\delta_{\mathbf{n}-\mathbf{e}_{j}},\;\mathbf{n}\in\mathbb{Z}^{d},\;j=1,\dots,d,$
@@ 2.1. Results
    with $\|\langle\mathbf{n},\mathbf{\alpha}\rangle\|_{d}=\mathrm{dist}(\langle\mathbf{n},\mathbf{\alpha}\rangle,\mathbb{Z})$ and the Fermi energy $\varepsilon_{F}\in\mathbb{R}$ ((cf. Theorem 2 ((i)) about the 11-d case below)).
@@
$(i)$ There is a one-dimensional version of Theorem 1 $(iv)$. It was considered in [13] where an explicit condition on the Cantor group $\Omega$ is formulated.
@@
    The one-dimensional Maryland model given by (2.1) – (2.3) for $d=1$, i.e. the one-dimensional discrete Schrödinger operator (2.13) with the potential ((cf. (2.5)))
@@
    and with the Fermi energy (1.6) $\varepsilon_{F}\in\sigma_{-}(H_{\omega}):=(-\infty,-\varepsilon_{0})$, where $\varepsilon_{0}\geq 0$ solves the equation
    (2.16) $\beta(\alpha)=\gamma(\varepsilon_{0},g),$
    with $\beta(\alpha)$ given in (2.14) and $\gamma(\lambda,g)\geq\gamma(0,g)>0$ being the Lyapunov exponent of the corresponding finite-difference equation of the second order (3.22) ((see (5.5) for the general definition of the Lyapunov exponent, and (3.23)–(3.25) for the form of $\gamma(\lambda,g)$ the one-dimensional Maryland model)). If $\gamma(0,g)\geq\beta(\alpha)$, then we have $\varepsilon_{F}\in\mathbb{R}$.
@@
The crucial ingredient of the proof of Theorem 1[(i)] and Theorem 2[(i)], is the following assertion of independent interest, according to which the eigenfunctions of the multi- and one-dimensional Maryland model are uniformly localised. This result provides a first example of uniformly localised eigenfunctions for an unbounded operator on the whole (infinite) spectrum. In the one-dimensional case the spectrum of the corresponding operator (the one-dimensional Schrödinger operator (2.13) with potential (2.15)) is described by similar formulae. However, since in this case the operator $W$ in (2.1) – (2.2) is the one-dimensional discrete Laplacian, an important part of the proof of the multidimensional estimate (2.20) can be made explicit. thereby leading to a stronger one-dimensional version.
Theorem 3.
  • (i)
    Let $d>1$, and let $H_{\omega}$ be the multidimensional Maryland model, the operator (2.1) – (2.2) with the potential (2.5), where $\alpha$ satisfies the multidimensional Diophantine condition (2.6). Then the spectrum of $H_{\omega}$ occupies the whole $\mathbb{R}$, is pure point and of multiplicity 1 ((see e.g. [41, Sections 18.A-18.B])). If $\{\lambda_{\mathbf{l}}(\omega)\}_{\mathbf{l}\in\mathbb{Z}^{d}}$ are its eigenvalues and $\{\psi_{\mathbf{l}}(\omega)\}_{\mathbf{l}\in\mathbb{Z}^{d}}$, $\psi_{\mathbf{l}}(\omega)=\{\psi_{\mathbf{l}}(\omega,\mathbf{n})\}_{\mathbf{n}\in\mathbb{Z}^{d}}$ are the corresponding eigenfunctions of $H_{\omega}$, then there exist $C<\infty$ and $c>0$ independent of $\lambda\in\mathbb{R}$ and of $\mathbf{n,l}\in\mathbb{Z}^{d}$ such that
    (2.20) $|\psi_{\mathbf{l}}(\lambda_{\mathbf{l}},\mathbf{n})|\leq Ce^{-c|\mathbf{n-l}|},\;\text{for any}\;\mathbf{n,l}\in\mathbb{Z}^{d}$
    i.e., the operators has uniformly localised eigenfunctions (2.24).
  • (ii)
    Let $H_{\omega}$ be the $1$-dimensional Maryland model (2.15), $\gamma(\lambda,g)\geq\gamma(0,g)>0$ be its Lyapunov exponent and $\varepsilon_{0}>0$ be the solution of the equation $\beta(\alpha)=\gamma(\varepsilon_{0},g)$ with $\beta(\alpha)$ of (2.14). Then the spectrum of $H_{\omega}$ is pure point on $\sigma_{\pm}({H_{\omega}}):=\{\pm\lambda\geq\varepsilon_{0}>0\}$, and if $\varepsilon_{0}=0$, then $\sigma_{pp}(H_{\omega})=\sigma_{-}(H_{\omega})\cup\sigma_{+}(H_{\omega})=\mathbb{R},$ see e.g. [41, Section 18.C].
    If $\{\lambda_{{l}}(\omega)\}_{{l}\in\mathbb{Z}}$ are its eigenvalues and $\{\psi_{{l}}(\omega)\}_{{l}\in\mathbb{Z}}$, $\psi_{{l}}(\omega)=\{\psi_{{l}}(\omega,{n})\}_{{n}\in\mathbb{Z}}$ are the corresponding eigenfunctions of $H_{\omega}$, then there exist $C<\infty$ and $c>0$ independent of $\lambda\in\mathbb{R}$ and of ${n,l}\in\mathbb{Z}$ such that
    with $\omega$-independent parameters $C\leq\infty,\;c>0$ and $\varepsilon_{0}\geq 0$, i.e., the operators has uniformly localised eigenfunctions (2.24).
The description of eigenfunctions of the Maryland model ((see (3.1) – (3.3))) provides the simplest and a quite explicit illustration of the notion of the localisation centers and the uniformly localised eigenfunctions. These spectral objects can be defined for any selfadjoint operator $H$ acting on $\ell^{2}(\mathbb{Z}^{d})$. Namely, $H$ is said to have uniformly localised eigenfunctions ((ULE)) if and only if $H$ has a complete set
@@
The proofs of Theorems 1-3 are given in Section 3.
Remark 2.3.
Using recent results of Jitomirskaya and Kachkovskiy [28, Corollary 5.14], it is possible to obtain a stronger version of Theorem 2 $(ii)$.
In our next result, we consider one-dimensional discrete Schrödinger operators (2.13) with potentials generated by a subshift of finite type (see Section 5), a very interesting class of hyperbolic dynamical systems exhibiting the so-called classical dynamical chaos, see e.g. [9]. For these ergodic operators, the spectral localisation outside of a discrete set of energies and for a range of coupling constants was recently proved in [6, 7].
@@
To prove Theorem 4 in Section 5 we have to substantially extend the existing spectral theory. We believe that this extension (see Theorem 5 below) is of independent interest for the spectral theory of this class of operators and completes the results of important recent works [7, 6].
@@
In these cases we define the potential by $V_{\omega}(n)=gv(T^{n}\omega)$ with $v:\Omega_{A}\rightarrow\mathbb{R}$ satisfying the assumptions of Theorem 4. Then, there exists $g_{0}>0$ such that for all $0<g\leq g_{0}$ Theorems 4 and 5 hold true. Moreover, as pointed out in [7, Remark 2.18], $g_{0}$ is not too small if $\alpha$ is not too small.
An important ingredient of the proof of Theorem 5 is the following lemma of independent interest. It provides a general condition for “strong” localisation, more precisely, for exponential dynamical localisation in expectation (2.33) for a bounded ergodic one-dimensional discrete Schrödinger operator n terms of the absence of quantum-mechanical tunneling between the distant domains of the system. Various similar conditions were often used in the analysis of Anderson localisation since the pioneering paper by Fröhlich and Spencer [20].
Let $M=\{M(j,k)\}_{j,k=-n}^{n}$ be a $(2n+1)\times(2n+1)$ matrix. Denote by
where $(M-\lambda)^{-1}(k,l)$ is the $(k,l)$ element of the matrix $(M-\lambda)^{-1}$, the set of the “badly”-controlled spectral parameters.
Lemma 2.5.
$Q_{I}(0,l)\leq 16e^{-\frac{c}{20}l}\equiv 16e^{-c^{\prime}l}.$
The proof of Lemma 2.5 is given in Section 4.
Remark 2.6.
Note that for the case of bounded one-dimensional Schrödinger operator with independent identically distributed potential ((aka Anderson model)) Lemma 2.5 combined with large deviation principle ((see [31] an referenes therein)) yields a short proof of the exponential dynamical localisation in expectation (2.33), the strongest form of Anderson localisation. The first proof of localisation for one-dimensional Anderson model is going back to [24], and [33]. Since then the problem has been attracting a considerable attention to this day, see e.g. [31] and references therein.
@@ 2.2. Comments and Related Results
The proof of Theorems 4 and 5 for the Schrödinger operators whose potential is generated by a subshift of finite type is using as an input the uniform positivity of the Lyapunov exponent and the uniform large deviation-type estimate from [6, 7]. In addition, we use special techniques of the theory of dynamical systems involving the Markov partitions and related Markov chains [1, 9].
@@ 3. Finite-difference operators with quasi-periodic potential
In this section, we present the proofs of Theorems 1 and 2 on the Area Law for several classes of non-random ergodic operators. We begin with the proof of Theorem 3, which is important as itself and crucial in the proofs of Theorems 1 and 2.
To the best of our knowledge, the only class of ergodic operators for which the Area Law is rigorously established (using Criterion 1) is that consisting of the Schrödinger operators with independent identically distributed random potentials (aka Anderson model) [18].
@@
that is, $\Omega=\mathbb{T}$ equipped with the Borel $\sigma$-algebra and $\mathbf{P}$ being the normalised Lebesgue measure.
@@
We start with the proof of Theorem 3 which we consider to be of independent interest.
3.1. Uniformly localised eigenfunctions for the $d$-dimensional, $d\geq 1$, Maryland model.
@@
Note first that the function $\lambda_{0}$ of (3.2) is the functional inverse of the Integrated Density of States $N(\lambda)$ of $H_{\omega}$ (see [41, Sections 4.B – 4.C] and [3, Sections 3.3 – 3.4] for its definition and properties). We have for the Maryland model
(3.4) $N(\lambda)=\int_{-\infty}^{\lambda}n(\lambda)d\lambda,$
i.e.,
(3.5) $\lambda_{0}\circ N=1,$
and
(3.6) $n(\lambda)=\frac{g}{(2\pi i)^{d}\pi}\int_{\mathbb{T}^{d}}|w(\mathbf{\eta})-\lambda+ig|^{-2}\prod_{j=1}^{d}\frac{d\eta_{j}}{\eta_{j}},$
is the Density of States of $H_{\omega}$ and $w:\mathbb{T}^{d}\rightarrow\mathbb{R}$ is given by the $d$-dimensional Fourier series with coefficients of (2.2)
(3.7) $\mathbf{\eta}$ $=(\eta_{1},\dots,\eta_{d})\in\mathbb{T}^{d},\,\,\mathbf{n}=(n_{1},\dots,n_{d})\in\mathbb{Z}^{d}.$
Next, the function $\psi_{0}$ of (3.1) is (see [41, Section 18.B])
(3.8) $\psi_{0}(\lambda,{\mathbf{n}})=\frac{1}{(2\pi i)^{d}(\pi n(\lambda)/g)^{1/2}}\int_{\mathbb{T}^{d}}\frac{e^{t(\lambda,\mathbf{\eta})}}{(w(\mathbf{\eta})-\lambda-ig)}\prod_{j=1}^{d}\eta^{n_{j}-1}d\eta_{j},$
where
(3.9) $t(\lambda,\mathbf{\eta})=\sum_{\mathbf{m}\in\mathbb{Z}^{d}}t_{\mathbf{m}}(\lambda)\prod_{j=1}^{d}\eta_{j}^{m_{j}},\;t_{\mathbf{m}}(\lambda)=\frac{l_{\mathbf{m}}(\lambda)}{(1-e^{i\langle\mathbf{\alpha,m}\rangle})}$
with
(3.10) $l_{\mathbf{m}}(\lambda)=(2\pi i)^{-d}\int_{\mathbb{T}^{d}}\log c(\lambda,\mathbf{\eta})\prod_{j=1}^{d}\eta^{m_{j}-1}d\eta_{j},$
and
(3.11) $c(\lambda,\mathbf{\eta})=-\frac{w(\mathbf{\eta})-\lambda+ig}{w(\mathbf{\eta})-\lambda-ig}.$
Thus, the $c(\lambda,\mathbf{\eta})$ of (3.11) is the input of the above analytic procedure given by formulae from (3.10) to (3.8) that leads to the eigenfunctions (3.3).
Our goal is to prove the bound (2.20). This combined with Remark 2.2 and Criterion 2 will allow us to establish the validity of the Area Law for the Maryland model.
We are going to use the following simple fact.
Lemma 3.1.
Let $f:\mathbb{T}^{d}\rightarrow\mathbb{C}$ be a periodic function and $\{F_{\mathbf{n}}\}_{\mathbf{n}\in\mathbb{Z}^{d}}$ be its Fourier coefficients. We have
where for any $0<\rho_{2}<\rho_{1}$ and $a=\rho_{1}-\rho_{2}>0$ it has the bound
[…]
(3.13) $F$ $=$ $\max_{|r|\leq\rho_{1},\;|\mathbf{\eta}|=1}|f(e^{r}\mathbf{\eta})|.$
Proof.
Now, applying Lemma 3.1 (i) to $w$ of (3.7) whose Fourier coefficients satisfy (2.2), we find that $w$ admits the analytic continuation to $\mathbb{T}_{\rho_{1}}^{d},\;0<\rho_{1}<\rho$, where $\rho>0$ is given in (2.2). Moreover, since the “initial” $w$ of (3.7) is real-valued ($w:\mathbb{T}^{d}\rightarrow\mathbb{R}$), we have for its analytic continuation
(3.14) $\lim_{r\rightarrow 0}\mathrm{Im}\ w(e^{r}\mathbf{\eta})=0,\;|\mathbf{\eta}|=1.$
(3.15) $|w(e^{r}\mathbf{\eta})-\lambda\pm ig|\geq|g\pm\mathrm{Im}\,w(e^{r}\mathbf{\eta})|\geq|g|-|\mathrm{Im}\,w(e^{r}\mathbf{\eta})|\geq g/2,$
where the last inequality follows from (3.14). We have then from (3.11) that $|\log c(\lambda,\mathbf{\eta})|\leq|\log|c(\lambda,\mathbf{\eta})|\ |+2\pi,$ and for any $|r|\leq\rho_{2}$
$\leq$ $2|g|\ |w(e^{r}\mathbf{\eta})-\lambda-ig|^{-1}\leq 4,$
where in obtaining the last bound of the r.h.s. above we used (3.15). Combining the two last bounds, we get for any $|r|\leq\rho_{2}$
$|\log c(\lambda,e^{r}\mathbf{\eta})|\leq C,$
where $C$ is a constant. Now combining (3.10) and Lemma 3.1 (ii) we obtain
$|l_{\mathbf{m}}(\lambda)|\leq Le^{-\rho_{2}|\mathbf{m}|},$
where $L<\infty$ and $\rho_{2}>0$ do not depend on $\lambda$ and $\mathbf{m}$.
This and the Diophantine condition (2.6) implies for the Fourier coefficients $t_{\mathbf{m}}(\lambda)$ of (3.9)
(3.16) $|t_{\mathbf{m}}(\lambda)|\leq Le^{-\rho_{2}^{\prime}|\mathbf{m}|},\;0<\rho^{\prime}_{2}<\rho_{2}.$
Then, Lemma 3.1 (ii) imply that $t(\lambda,\cdot)$ of (3.9) admits the analytic continuation into $\mathbb{T}_{\rho_{3}}^{d}$ with some $0<\rho_{3}<\rho^{\prime}_{2}$ and is bounded there for any $|r|\leq\rho_{3}$, i.e.,
(3.17) $|t(\lambda,e^{r}\mathbf{\eta})|\leq T<\infty,\;|r|\leq\rho_{3},$
where $T$ and $\rho_{3}$ are $\lambda$-independent according to (3.12).
Next, (3.15), and the analyticity of $w$ in $\mathbb{T}_{\rho_{1}}^{d}$ imply that
is analytic in $\mathbf{z}$ in $\mathbb{T}_{\rho_{3}}^{d}$ for any $\lambda\in\mathbb{R}$. Since the r.h.s. of (3.8) is the Fourier coefficient of $h(\lambda,\mathbf{\cdot})$, (3.13) yields the following bound for the r.h.s. of (3.8)
(3.18) $\Psi e^{-\rho_{3}|\mathbf{n}|},\;\mathbf{n}\in\mathbb{Z}^{d},\;\Psi=\max_{\lambda\in\mathbb{R},\;|r|\leq\rho_{3},\;|\mathbf{\eta}|=1}|h(\lambda,e^{r}\mathbf{\eta})|.$
It follows from (3.6) that
  1. (1)
    $(n(\lambda))^{-1/2}\geq(\pi g)^{1/2},$ and $(n(\lambda))^{1/2}=|\lambda|^{-1}(1+o(1)),\;|\lambda|\rightarrow\infty$, by (3.6);
  2. (2)
    $|w(\mathbf{z})-\lambda-ig|,\;|\mathbf{z}|\in[e^{-\rho_{3}},e^{\rho_{3}}]$ is bounded from below by (3.15) and is $|\lambda|(1+o(1)),\;|\lambda|\rightarrow\infty$;
  3. (3)
    $e^{t(\lambda,\mathbf{z})}\leq e^{T}$ by (3.17).
This and (3.18) imply that $\Psi$ is independent of $\lambda\in\mathbb{R}$ and $\mathbf{n}\in\mathbb{Z}^{d}$.
Thus we obtain
@@
with $\Psi$ of (3.18) as $C$ and $\rho_{3}$ of (3.17) as $c$. Combining (3.3) and (3.19), we obtain
@@
where $C<\infty$ and $c>0$ are independent of $\omega$. ∎
Proof of Theorem 3 $(ii)$.
According to [41, Section 18.C], the pure point spectrum consists of two semi-infinite components (2.16)
Since in this case operator $W$ in (2.1) is the one-dimensional discrete Laplacian, we set $W_{0}=0,\;W_{\pm 1}=1$ in (2.2). Hence, $w$ of (3.7) is
(3.21) $w(\eta)=\eta+\eta^{-1}.$
Thus, $w$ is analytic everywhere except zero and infinity, i.e., in this case the analog of $\rho_{1}$ of the proof above is infinity.
Next, the integral in the r.h.s. of formula (3.10) for $d=1$ and $w$ of (3.21) can be calculated yielding (cf. (3.10))
$l_{m}=\frac{2i}{|m|}e^{-\gamma|m|}\sin|m|\varphi.$
(3.22) $u_{n+1}+u_{n-1}+g\tan(\alpha n+\omega)u_{n}=\lambda u_{n}.$
We have
(3.23) $\mathrm{sinh}\gamma(\lambda,g)=s>0,$
where $s>0$ is such that
(3.24) $s^{4}+(4-\lambda^{2}-g^{2})s^{2}-g^{2}=0,$
(3.25) $\gamma(\lambda,g)\geq\gamma(0,g)=\mathrm{arc\,sinh}(|g|/2)>0,\;g\neq 0.$
$\rho^{\prime}:=\gamma(\varepsilon_{0},g)-\beta(\alpha)>0,$
we obtain from (3.9) and (3.10) for the Fourier coefficients $t_{m}(\lambda)$ of function $t(\mathbf{\eta},\cdot)$ of (3.9) for $d=1$
This bound is an analog of (3.16). Hence, repeating the argument of the previous proof starting from (3.9) we get
with $\omega$-independent parameters $C<\infty,\;c>0$ and $\varepsilon_{0}\geq 0$. ∎
3.2. Proof of Theorem 1
Proof of Theorem 1 $(i)$.
Spectral theorem for the Fermi projection